Calculate the pH of a mixture containing 0.01 M NaCH3CO2 and 0.02 M CH3CO2H given the pKa of CH3CO2H = 4.75


1
Expert's answer
2020-12-01T09:12:11-0500

Given a solution which is 0.01M sodium acetate and 0.02M acetic acid.This is an acidic buffer.

And, pKa=4.75

Hence, pH=pKa+log(CH3COONa)/(CH3COOH)

=4.75+log(0.01/0.02)

=4.75-0.301

=4.449

So pH of the solution is 4.449